# Steady state diagram of interacting fermionic atoms coupled to dissipative cavities

### v1

@@ V.1.1 Self-organization phase transition
Figure 5: Cuts of the (a), (c) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b), (d) inverse temperature $\beta$ as a function of atoms-cavity coupling $\hbar g/J$ of a finite size system L=8L\!=\!8 (blue points) and of the weak tunneling perturbation in the thermodynamic limit (green points) at quarter filling. The parameters used are $\hbar\delta/J\!=\!8$ and $\hbar\varGamma/J\!=\!1$ and in (a), (b) at U/J=8U/J\!=\!8 and in (c), (d) at U/J=20U/J\!=\!20. Vertical lines denote the critical coupling $g_{\text{cr}}/J$ and resonances. The orange dashed line in (a) is the approximate scaling for $U/J\!\to\!\infty$ [Eqs. (40)].
@@
Figure 6: Simplified phase diagrams varying the atoms-cavity coupling $\hbar g/J$ and (a) the on-site interaction $U/J$ at $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$; (b) the pump-cavity detuning $\hbar\delta/J$ at U/J=8U/J\!=\!8, $\hbar\varGamma/J\!=\!1$; (c) the dissipation rate $\hbar\varGamma/J$ at $\hbar\delta/J\!=\!6$, U/J=8U/J\!=\!8 of a finite size system L=8L\!=\!8 at quarter filling. The green line marks the transition between normal (purple region) and self-organized phase (blue region) determined by solving Eqs. (5) & (15) with ED. The red dashed lines in (b), (c) is calculated from Eq. (36) with Eq. (37d) using the self-consistently determined $\beta$ in the limit $U/J\!\to\!\infty$. For the purple dashed lines each occurrence of $\beta$ is replaced by the large temperature limit Eq. (47) before solving. The blue dashed line in (a) is the equivalent result for U/J=0U/J\!=\!0. The solid black line marks the critical coupling $\hbar g^{\text{MF}}_{\text{cr}}$ obtained for the T=0T\!=\!0 MF approach solving Eq. (5) via ED. We compare to the black dashed line, the approximate analytical result for $g_{\text{cr}}^{\text{MF}}$ [Eq. (36) with Eq. (37b)]. The small hatched triangular region on the center right marks the fluctuation-induced bistability region within the self-organized phase.
We begin by analyzing the self-organization transition of the atoms-cavity system. In Fig. 5 we show both the T=0T\!=\!0 MF results (black line) and also the results obtained taking the fluctuations into account (many-body adiabatic elimination method derived in Sec. III.2) for two different interaction strength $U/J\!=8~\text{and}~20$, respectively. The blue/red symbols are the results for exact diagonalization and green symbols are from the weak tunneling perturbation in the thermodynamic limit (Sec. III.4). At low atoms-cavity coupling strength $g$, the system is in a normal state characterized by a vanishing photon number and vanishing atomic sublattice density imbalance [Fig. 5 (a),(c)]. As the atoms-cavity coupling $g$ is increased above $g_{\text{cr}}$ the system enters into the self-ordered phase, signaled by a simultaneous increase of the cavity field (not shown) and the sublattice imbalance, see for example the behavior of $\langle\hat{\Delta}\rangle/L^{d}$ around $\hbar g_{\text{cr}}/J\!\approx\!3$ for the T=0T\!=\!0 MF result and around $\hbar g_{\text{cr}}/J\!\approx\!5.3$ taking the fluctuations into account for the parameters considered in Fig. 5 (a). In both approaches, the self-ordering transition is present. However, the exact location and the nature of the transition can be very distinct as we will discuss in more detail below.
@@
To understand better the observed behavior, we derive expressions for the critical coupling in the solvable non-interacting limit (U/J=0U/J\!=\!0) and the $U/J\!\to\!\infty$ limit. In the $U/J\!\to\!\infty$ limit this is done by approximating the fermions by effectively spinless fermions. Similar approaches at the T=0T\!=\!0 MF level were derived for non-interacting semiclassical particles without external optical lattice Asbóth et al. (2005); Niedenzu et al. (2011) as well as bosons Kirton et al. (2019); Nagy et al. (2008). To compute an approximate expression for the critical coupling $g_{\text{cr}}$, one has to investigate for which value of the coupling the implicit equation can be solved to obtain a small non-zero value of $\lambda$. The analytical expression for $g_{\text{cr}}$ derived are thus valid for incommensurate values of the filling, where one can assume the cavity field, $\lambda$, to be small close to the phase transition within the self-ordered regime. The derivation is given in appendix A for a one-dimensional system. However, the equivalent expressions for higher dimensions can be obtained in an analogous way. We obtain an expression for the critical coupling given by
@@
The function F depends on the approximation used. For the T=0T\!=\!0 MF approach, the function F only depends on the filling of the atomic system. In the thermodynamic limit $F\!=\!\text{F}\!\left(\left.n\pi/2\right|1\right)$ where
@@
For the finite system the momentum sum has to be taken over all occupied momenta (black dashed line in Fig. 6). When considering $U/J\!\to\!\infty$, the ground state is approximated by spinless fermions filling the lower band completely (gray dashed line) at quarter filling. Close to the transition we further obtain the scaling of $\lambda$ (see appendix A) given by
@@
This scaling agrees with the one derived for the Dicke model $\alpha\!\propto\!(g^{\prime}-g^{\prime}_{\text{cr}})^{1/2}$ Kirton et al. (2019) with $g^{\prime}$ the respective coupling strength and $\alpha\!=\!\langle\hat{a}\rangle$ the cavity order parameter. The scaling exponent $1/2$ is obtained both with and without dissipation. For the finite temperature approach the function becomes more involved and is given by
@@
which depends on the self-consistently evaluated $\beta$ (see appendix A). Therefore, the function F in this case can also depend on the system parameters [see red dashed line in Fig. 6 (b),(c)]. The self-ordering transition of an ensemble of non-interacting semiclassical particles Asbóth et al. (2005); Niedenzu et al. (2011) shows the same scaling in a mean-field approach. However, in Ginzburg-Landau type theories a switch in the scaling to $1/4$ at a tricritical point is derived for low pump-cavity detuning Keeling et al. (2014).
@@
The obtained critical coupling is shown as a purple dashed line in Fig. 6. Thus, including the self-consistent effective temperature leads to a different critical coupling, the scaling of which we will discuss below. A similar result was obtained for bosonic atoms Bezvershenko et al. (2021).
As expected in both approaches, the non-interacting limit agrees perfectly with the numerically determined critical coupling [see Fig. 6 (a)], if the finite system expression for F is taken [see Eq. (37b) or Eq. (37d)]. Small deviations are seen for the $U/J\!\to\!\infty$ limit, since here the assumption to use spinless fermions is only an approximation.
@@ V.1.2 Fluctuation-effects in the ordered phase
As in our results the excited states of the effective model $\hat{H}_{\text{eff}}$ in the steady state density matrix are present, it is important to understand the nature of the excited states and the behavior of the gaps. We investigate the excitation gap from the ground state in the energy spectrum $E_{\text{gap}}$ that is obtained from the eigenspectrum of $\hat{H}_{\text{eff}}$. At low filling, in the atomic limit we obtain an approximate band structure where all eigenstates in the lowest energy band are defined by maximal sublattice density imbalance $\langle\hat{\Delta}\rangle\!=\!N$ and exclusively singly occupied sites $\sum_{j}\langle\hat{n}_{j\uparrow}\hat{n}_{j\downarrow}\rangle\!=\!0$. We calculate the excitation gap between the lowest energy band and the lowest energy eigenstate with differing $\langle\hat{\Delta}\rangle$ or $\sum_{j}\langle\hat{n}_{j\uparrow}\hat{n}_{j\downarrow}\rangle$. In the parameter regime considered here the finite gap between states with equal density imbalance and number of double occupancies, occurring from the lifting of the degeneracy by the finite size effects and higher order hoppings, can be omitted. As shown in Fig. 7, the excitation gap (green curve) increases above the phase transition and approaches $U/J$ for $2\hbar g\lambda\!\gg\!U$. Thus, the nature of the excitation gap changes from band-insulator to Mott-insulator on the low-potential sublattice sites.
@@
Figure 8: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta$ as a function of atoms-cavity coupling $\hbar g/J$ and on-site interaction $U/J$ of a finite size system L=8L\!=\!8 at quarter filling. The parameters used are $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$. The small hatched triangular region on the center right marks the fluctuation-induced bistability region within the self-organized phase. Lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances.
By increasing the coupling strength above the $g_{\text{cr}}$ we observe that the sublattice imbalance $\langle\hat{\Delta}\rangle$ and the effective temperature have a non-monotonic behavior. The imbalance shows an increase above $g_{\text{cr}}$ to a maximum between $\hbar g/J\!\approx\!5$ to $6$ in Fig. 5 (a),(c). At even higher coupling strength one or two further maxima are seen (which will be analyzed further in Sec. V.1.3), followed by a decrease for larger couplings. Also the inverse temperature $\beta$ shows a strong increase to a maximum followed by one or two additional maxima [Fig. 5 (b),(d)] on an overall decrease. These maxima corresponds to a cooling effects, and a local minimum in temperature on the overall heating. The atoms are cooled by the cavity in this regime. We find that the first minimum in temperature occurs approximately when the photon energy matches the height of the self-organized potential, $2g\lambda\!\approx\!\delta$ (purple line in Fig. 5). The position of the first maximum in the imbalance and the corresponding minimum in the effective temperature can be very well followed in Fig. 8 and Fig. 9, where we show the dependence on the different system parameters. In both figures the first maximum/minimum for the imbalance/inverse temperature follows approximately the purple line indicating the condition $2g\lambda\!\approx\!\delta$.
@@
Figure 9: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L$ and (b) inverse temperature $\beta$ as a function of atoms-cavity coupling $\hbar g/J$ and detuning $\hbar\delta/J$ of a finite size system L=8L\!=\!8 at quarter filling. The parameters used are U/J=8U/J\!=\!8 and $\hbar\varGamma/J\!=\!1$. Lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances in the atomic limit.
Figure 10: Sketches of the energy transfer transitions captured by Eq. (29a). (a) First line in Eq. (29a): hopping from a high- to an empty neighboring low potential sublattice site, adding a photon to/removing a photon from the cavity field $\Delta E_{nm}\!=\!\pm(2\hbar g\lambda\!\pm\!\hbar\delta)$. (b) Second line in Eq. (29a) with p=1p\!=\!1: hopping from a high- to a neighbouring low potential sublattice site, creating a double occupancy and adding/removing a photon to the cavity field $\Delta E_{nm}\!=\!\pm(2\hbar g\lambda\!-\!U\!\pm\!\hbar\delta)$. (c) Next neighbour hopping process resonant at $\Delta E_{nm}\!=\!\pm(U\!-\!\hbar\delta)$
@@
Figure 11: The (a) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta$ as a function of $\hbar\varGamma/J$ of a finite size system L=8L\!=\!8 at quarter filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid line. The parameters are $\hbar g/J\!=\!5$, U/J=8U/J\!=\!8 and $\hbar\delta/J\!=\!2$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances. The orange dashed line is the approximate scaling for $U/J\!\to\!\infty$ [Eqs. (40)].
@@
Figure 12: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta$ as a function of atoms-cavity coupling $\hbar g/J$ and dissipation $\hbar\varGamma/J$ of a finite size system L=8L\!=\!8 at quarter filling. The parameters used are U/J=8U/J\!=\!8, $\hbar\delta/J\!=\!6$. Lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances in the atomic limit.
@@
Figure 13: The (a) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta$ as a function of $U/J$ of a finite size system L=8L\!=\!8 at quarter filling. The T=0T\!=\!0 MF solution is shown as black solid line in (a). The fixed parameters used are $\hbar g/J\!=\!8$, $\hbar\delta/J\!=\!8$ and $\hbar\varGamma/J\!=\!1$. Vertical lines denote resonances, in the atomic limit.
@@
Figure 14: The (a) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta$ as a function of $\hbar\delta/J$; of a finite size system L=8L\!=\!8 stable (blue); in the thermodynamic limit at low hopping [Eqs. (29)] stable (dark green) and unstable (light green) at quarter filling. The T=0T\!=\!0 MF solution is shown as black solid line. The parameters are $\hbar g/J\!=\!6$, U/J=8U/J\!=\!8 and $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances.
@@
For $U/J\!\to\!\infty$ and in the strong coupling regime, we have
@@
We note that the result for spinful fermions obtained here differs from the one obtained for a Bose-Hubbard system in Ref. Bezvershenko et al. (2021), where the photon number scales as n(n+1)n(n\!+\!1) with the particle density. We observe in our results both the scaling of the temperature, $\beta\!\propto\!g^{-2}$, and the sublattice imbalance, $\langle\hat{\Delta}\rangle\!\propto\!g^{-1}$, as we increase the coupling strength (see orange dashed line in Fig. 5). The obtained dependence on $\varGamma$ also matches the results (see orange dashed line in Fig. 11). The expressions in Eq. (39) and Eq. (40) are in stark contrast with the expectations from the T=0T\!=\!0 MF approach.
@@
Figure 15: Cuts of the (a) scaled photon number, (b) inverse temperature $\beta$, (c) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (d) average double-occupancy as a function of atoms-cavity coupling $\hbar g/J$ at U/J=40U/J\!=\!40; of a finite size system L=8L\!=\!8 stable (blue) and unstable (red); in the thermodynamic limit at low hopping [Eqs. (29)] stable (dark green) and unstable (light green) at quarter filling. The T=0T\!=\!0 MF solution is shown as black solid line. The other parameters used are $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances. The orange dashed lines are the scalings of the analytical approximations for $U/J\!\to\!\infty$ [Eqs. (40)].

### v2

@@ V.1.1 Self-organization phase transition
Figure 5: Cuts of the (a), (c) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b), (d) inverse temperature $\beta J$ as a function of atoms-cavity coupling $\hbar g/J$ of a finite size system L=8L\!=\!8 for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid line. We also show stable (dark green) and unstable (light green) results in the weak tunneling perturbation in the thermodynamic limit at quarter filling. The parameters used are $\hbar\delta/J\!=\!8$ and $\hbar\varGamma/J\!=\!1$ and in (a), (b) at U/J=8U/J\!=\!8 and in (c), (d) at U/J=20U/J\!=\!20. Vertical lines denote the critical coupling $g_{\text{cr}}/J$ and resonances as explained in the legend. The orange dashed line in (a) is the approximate scaling for $U/J\!\to\!\infty$ [Eqs. (40)].
@@
Figure 6: Simplified phase diagrams varying the atoms-cavity coupling $\hbar g/J$ and (a) the on-site interaction $U/J$ at $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$; (b) the pump-cavity detuning $\hbar\delta/J$ at U/J=8U/J\!=\!8, $\hbar\varGamma/J\!=\!1$; (c) the dissipation rate $\hbar\varGamma/J$ at $\hbar\delta/J\!=\!6$, U/J=8U/J\!=\!8, for a finite size system L=8L\!=\!8 at quarter filling. The green line marks the transition between normal (purple region) and self-organized phase (blue region) at $\hbar g_{\text{cr}}/J$ determined by solving Eqs. (5) & (15) with ED. The red dashed line in (b), (c) is calculated from Eq. (36) with Eq. (37d) using the self-consistently determined $\beta$ in the limit $U/J\!\to\!\infty$. For the purple dashed lines the large temperature limit Eq. (38) is used. The blue dashed line in (a) is the equivalent result for U/J=0U/J\!=\!0. The solid black line marks the critical coupling $\hbar g^{\text{MF}}_{\text{cr}}$ obtained for the T=0T\!=\!0 MF approach solving Eq. (5) via ED. We compare to the black (gray) dashed line, the approximate analytical result for $g_{\text{cr}}^{\text{MF}}$ [Eq. (36) with Eq. (37b)] for U/J=0U/J\!=\!0 ($U/J\!\to\!\infty$). The small hatched triangular region on the center right marks the fluctuation-induced bistability region within the self-organized phase.
We begin by analyzing the self-organization transition of the atoms-cavity system. In Fig. 5 we show both the T=0T\!=\!0 MF results (black line) and also the results obtained taking the fluctuations into account (many-body adiabatic elimination method derived in Sec. III.2) for two different interaction strength $U/J\!=8~\text{and}~20$, respectively. The blue/red symbols are the results for exact diagonalization and green symbols are from the weak tunneling perturbation in the thermodynamic limit (Sec. III.4). At low atoms-cavity coupling strength $g$, the system is in a normal state characterized by a vanishing photon number and vanishing atomic sublattice density imbalance [Fig. 5 (a),(c)]. As the atoms-cavity coupling $g$ is increased above $g_{\text{cr}}$ the system enters into the self-ordered phase, signaled by a simultaneous increase of the cavity field (not shown) and the sublattice imbalance, see for example the behavior of $\langle\hat{\Delta}\rangle/L^{d}$ around $\hbar g_{\text{cr}}/J\!\approx\!3$ for the T=0T\!=\!0 MF result and around $\hbar g_{\text{cr}}/J\!\approx\!5.3$ taking the fluctuations into account for the parameters considered in Fig. 5 (a). In both approaches, the self-ordering transition is present. However, the exact location and the nature of the transition can be very distinct as we discuss in more detail below.
@@
To understand better the observed behavior, we derive analytical expressions for the critical coupling in the solvable non-interacting limit (U/J=0U/J\!=\!0). Similar approaches at the T=0T\!=\!0 MF level were derived for non-interacting semiclassical particles without external optical lattice Asbóth et al. (2005); Niedenzu et al. (2011) as well as bosons Kirton et al. (2019); Nagy et al. (2008). Assuming that $\lambda$ is continuous at the transition, one can compute an approximate expression for the critical coupling $g_{\text{cr}}$, by searching for which value of the coupling the implicit equation can be solved to obtain a small non-zero value of $\lambda$ while using expansions for small $\lambda$. The assumption of a continuous $\lambda$ is valid for incommensurate values of the filling. The derivation of the critical coupling $g_{\text{cr}}$ is given in Appendix A for a one-dimensional system. The obtained expression is given by
@@
The equivalent expressions for higher dimensions can be obtained in an analogous way. The function F depends on the approximation used. We first describe in Eq. (37a) and Eq. (37b) the results obtained for the T=0T\!=\!0 MF approach and in Eq. (37d) the results obtained with the approach including fluctuations around the mean-field coupling. For the T=0T\!=\!0 MF approach the function F only depends on the filling of the atomic system. In the thermodynamic limit $F\!=\!\text{F}^{\text{MF},L\to\infty}\!\left(\left.n\pi/2\right|1\right)$ where
@@
For the finite system the momentum sum has to be taken over all occupied momenta [black dashed line in Fig. 6 (a)]. This approximation agrees very well with the numerical results obtained at low interaction strength using the T=0T\!=\!0 MF approximation (black line in Fig. 6).
We can make use of the derived expression to also obtain an approximation at $U/J\!\to\!\infty$. Here we assume that due to the infinite repulsion the ground state is approximated by non-interacting, but spinless, fermions with double the filling. For example at quarter filling, the spinless fermions would completely fill the lower band [gray dashed line in Fig. 6 (a)]. The numerical results at large interaction strength only slightly deviate from the results obtained within this approximation.
Close to the transition we further obtain the scaling of $\lambda$ (see Appendix A) given by
@@
This scaling agrees with the one derived for the Dicke model $\alpha\!\propto\!(g^{\prime}-g^{\prime}_{\text{cr}})^{1/2}$ Kirton et al. (2019)
with $g^{\prime}$ the respective coupling strength and $\alpha\!=\!\langle\hat{a}\rangle$ the cavity order parameter. The mean-field Dicke scaling exponent $1/2$ is the same with and without photon losses. The self-ordering transition of an ensemble of non-interacting semiclassical particles Asbóth et al. (2005); Niedenzu et al. (2011) shows the same scaling in a mean-field approach. However, in Ginzburg-Landau type theories a change in the scaling to $1/4$ at a tricritical point is derived for low pump-cavity detuning Keeling et al. (2014).
For the finite temperature approach the function $F$ becomes more involved and is given by
@@
which depends on the self-consistently evaluated $\beta$ (see Appendix A). Therefore, the function F in this case cannot be easily approximated by an analytical expression. Its evaluation is shown as the red dashed line in Fig. 6 (b)-(c) for the infinitely strongly interacting gas.
@@
The obtained critical coupling is shown as blue dashed line in Fig. 6 (a) for the non-interacting gas and as the purple dashed line in Fig. 6 (a)-(c) for the infinitely strongly interacting gas, using the spinless fermion approximation described above.
Thus, including the self-consistent effective temperature leads to a different critical coupling, the scaling of which we will discuss below. A similar result was obtained for bosonic atoms Bezvershenko et al. (2021).
@@ V.1.2 Fluctuation-effects in the ordered phase
As in our results the excited states of the effective model $\hat{H}_{\text{eff}}$ are present in the steady state density matrix, it is important to understand the nature of the excited states and the behavior of the gaps. We investigate the excitation gap from the ground state in the energy spectrum $E_{\text{gap}}$ that is obtained from the eigenspectrum of $\hat{H}_{\text{eff}}$. At low filling, in the atomic limit we obtain an approximate band structure where all eigenstates in the lowest energy band are defined by maximal sublattice density imbalance $\langle\hat{\Delta}\rangle\!=\!N$ and exclusively singly occupied sites $\sum_{j}\langle\hat{n}_{j\uparrow}\hat{n}_{j\downarrow}\rangle\!=\!0$. We calculate the excitation gap between the lowest energy band and the lowest energy eigenstate with differing $\langle\hat{\Delta}\rangle$ or $\sum_{j}\langle\hat{n}_{j\uparrow}\hat{n}_{j\downarrow}\rangle$. In the parameter regime considered here the finite gap between states with equal density imbalance and number of double occupancies, occurring from the lifting of the degeneracy by the finite size effects and higher order hoppings, can be omitted. As shown in Fig. 7, the excitation gap (green curve) increases above the phase transition and approaches $U/J$ for $2\hbar g\lambda\!\gg\!U$. Thus, the nature of the excitation gap changes from band-insulator to Mott-insulator on the low-potential sublattice sites.
@@
Figure 8: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta J$ as a function of atoms-cavity coupling $\hbar g/J$ and on-site interaction $U/J$ of a finite size system L=8L\!=\!8 at quarter filling. The parameters used are $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$. The small hatched triangular region on the center right marks the fluctuation-induced bistability region within the self-organized phase. Lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances as explained in the legend.
By increasing the coupling strength above the $g_{\text{cr}}$ we observe that the sublattice imbalance $\langle\hat{\Delta}\rangle$ and the effective temperature have a non-monotonic behavior. The imbalance shows an increase above $g_{\text{cr}}$ to a maximum between $\hbar g/J\!\approx\!5$ to $6$ in Fig. 5 (a),(c). At even higher coupling strength one or two further maxima are seen (which will be analyzed further in Sec. V.1.3), followed by a decrease for larger couplings. Also the inverse temperature $\beta$ shows a strong increase to a maximum followed by one or two additional maxima [Fig. 5 (b),(d)] on an overall decrease. These maxima corresponds to cooling effects, and a local minimum in temperature on the overall heating. The atoms are cooled by the cavity in this regime. We find that the first minimum in temperature occurs approximately when the photon energy matches the height of the self-organized potential, $2g\lambda\!\approx\!\delta$ (purple line in Fig. 5). The position of the first maximum in the imbalance and the corresponding minimum in the effective temperature can be very well followed in Fig. 8 and Fig. 9, where we show the dependence on the different system parameters. In both figures the first maximum/minimum for the imbalance/inverse temperature follows approximately the purple line indicating the condition $2g\lambda\!\approx\!\delta$.
@@
Figure 9: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L$ and (b) inverse temperature $\beta J$ as a function of atoms-cavity coupling $\hbar g/J$ and detuning $\hbar\delta/J$ of a finite size system L=8L\!=\!8 at quarter filling. The parameters used are U/J=8U/J\!=\!8 and $\hbar\varGamma/J\!=\!1$. Lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances in the atomic limit.
Figure 10: Sketches of the energy transfer transitions captured by Eq. (29a). (a) First line in Eq. (29a): hopping from a high- to an empty neighboring low potential sublattice site, adding a photon to/removing a photon from the cavity field $\Delta E_{nm}\!=\!\pm(2\hbar g\lambda\!\pm\!\hbar\delta)$. (b) Second line in Eq. (29a) with p=1p\!=\!1: hopping from a high- to a neighboring low potential sublattice site, creating a double occupancy and adding/removing a photon to the cavity field $\Delta E_{nm}\!=\!\pm(2\hbar g\lambda\!-\!U\!\pm\!\hbar\delta)$. (c) Next neighbor hopping process resonant at $\Delta E_{nm}\!=\!\pm(U\!-\!\hbar\delta)$
@@
Figure 11: The (a) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta J$ as a function of $\hbar\varGamma/J$ of a finite size system L=8L\!=\!8 at quarter filling for stable (blue) and unstable (red) results. The T=0T\!=\!0 MF solution is shown as black solid line. The parameters are $\hbar g/J\!=\!5$, U/J=8U/J\!=\!8 and $\hbar\delta/J\!=\!2$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances. The orange dashed line is the approximate scaling for $U/J\!\to\!\infty$ [Eqs. (40)].
@@
Figure 12: (a) Sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta J$ as a function of atoms-cavity coupling $\hbar g/J$ and dissipation $\hbar\varGamma/J$ of a finite size system L=8L\!=\!8 at quarter filling. The parameters used are U/J=8U/J\!=\!8, $\hbar\delta/J\!=\!6$. Lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances in the atomic limit.
@@
Figure 13: The (a) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta J$ as a function of $U/J$ of a finite size system L=8L\!=\!8 at quarter filling. The T=0T\!=\!0 MF solution is shown as black solid line in (a). The fixed parameters used are $\hbar g/J\!=\!8$, $\hbar\delta/J\!=\!8$ and $\hbar\varGamma/J\!=\!1$. Vertical lines denote resonances, in the atomic limit.
@@
Figure 14: The (a) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (b) inverse temperature $\beta J$ as a function of $\hbar\delta/J$; of a finite size system L=8L\!=\!8 stable (blue); in the weak tunneling perturbation in the thermodynamic limit [Eqs. (29)] stable (dark green) and unstable (light green) at quarter filling. The T=0T\!=\!0 MF solution is shown as black solid line. The parameters are $\hbar g/J\!=\!6$, U/J=8U/J\!=\!8 and $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances.
@@
While for $U/J\!\to\!\infty$ in the strong coupling regime we have
@@
We note that the result for spinful fermions obtained here differs from the one obtained for a Bose-Hubbard system in Ref. Bezvershenko et al. (2021), where the photon number scales as n(n+1)n(n\!+\!1) with the particle density. We observe in our results both the scaling of the temperature, $\beta\!\propto\!g^{-2}$, and the sublattice imbalance, $\langle\hat{\Delta}\rangle\!\propto\!g^{-1}$, as we increase the coupling strength (see orange dashed line in Fig. 5). The obtained dependence on $\varGamma$ also matches the results (see orange dashed line in Fig. 11). In Fig. 11 and Fig. 12 one also sees clearly that for large values of the dissipation heating is induced in the atomic system. With increasing dissipation the effective temperature rises approximately as $k_{B}T\sim\varGamma^{2}$ as predicted in Eq. (21). Only the resonances described in Sec. V.1.3 lead to a strongly non-monotonic behavior.
The expressions in Eq. (39) and Eq. (40) are in stark contrast with the expectations from the T=0T\!=\!0 MF approach.
@@
Figure 15: Cuts of the (a) scaled photon number, (b) inverse temperature $\beta J$, (c) sublattice imbalance $\langle\hat{\Delta}\rangle/L^{d}$ and (d) average double-occupancy as a function of atoms-cavity coupling $\hbar g/J$ at U/J=40U/J\!=\!40; of a finite size system L=8L\!=\!8 stable (blue) and unstable (red). The light blue interval marks the bistability region; in the weak tunneling perturbation in the thermodynamic limit [Eqs. (29)] stable (dark green) and unstable (light green) at quarter filling. The T=0T\!=\!0 MF solution is shown as black solid line. The other parameters used are $\hbar\delta/J\!=\!8$, $\hbar\varGamma/J\!=\!1$. Vertical lines denote the critical coupling $\hbar g_{\text{cr}}/J$ and resonances. The orange dashed lines are the scalings of the analytical approximations for $U/J\!\to\!\infty$ [Eqs. (40)].